Your calculated pressure should be in units of atm. Do not enter the units in the answer box below. A gas cylinder (container with a fixed volume) of nitrogen (N2) had a pressure of 4.64 atm at a temperature of 24.9 ºC. A fire in the lab raised the temperature to 278.7 ºC. What is new the pressure in the cylinder in atm?
Answer #1

Given that: Constant volume.

Hence, Gay-Lussac's law can be applied:

P1/T1 = P2/T2

We have:

P1 = 4.64 atm

T1 = 24.9oC = 24.9 + 273.15 K = 298.05 K

T2 = 278.7oC = 278.7 + 273.15 K = 551.85 K

Therefore, Final pressure will be:

P1/T1 = P2/T2

P2 = (P1/T1) x (T2)

P2 = (4.64 / 298.05) x ( 551.85 K)

P2 = 8.59 atm ------------------------ (**Answer**)